# A new machine learning framework for occupational accidents forecasting with safety inspections integration

### v1

@@ 3.1 Forecasting accident risk via binary time series modeling
We represent the daily occurrence of accidents using a binary time series $\{y_{t}\}_{t=1}^{T}$, where $y_{t}=1$ if at least one accident occurs on day $t$ and $y_{t}=0$ otherwise. Two predictor families are distinguished: (i) static calendar covariates $s_{t}$ such as month or day of week, which are fully known for any future date; and (ii) dynamic inspection covariates $c_{t}$ extracted from the most recent safety inspection report available at day $t$.
@@
3.2 Multi-step forecasting strategies
Our aim is to predict the sequence $\{y_{t+h}\}_{h=1}^{H}$, thus producing forecasts for multiple future time steps. Several strategies can be adopted (Bontempi et al., 2013), which are commonly categorised by the output dimensionality of the underlying model.
@@ 3.2.1 Direct recursive strategy (DirRec)
A single-output learner such as logistic regression (Bishop and Nasrabadi, 2006), random forests (Breiman, 2001), or gradient-boosted trees like XGBoost (Chen and Guestrin, 2016) yields a one-dimensional output. To handle multi-step forecasts, we use the DirRec (Direct-Recursive) strategy (Sorjamaa and Lendasse, 2006), which combines direct and recursive methods. It trains a separate estimator $f_{h}(\cdot;\theta_{h})$ for every horizon $h=1,\dots,H$. Except for the first, each estimator receives the forecasts produced at earlier horizons as additional inputs. Accordingly, the one-step-ahead forecast is
$\hat{p}_{t+1}\;=\;f_{1}\!\bigl{(}y_{t},\dots,y_{t-d_{y}+1},\;C_{t}^{d_{c}},\;s_{t+1};\theta_{1}\bigr{)},$
@@
$\hat{p}_{t+h}\;=\;f_{h}\!\bigl{(}\hat{p}_{t+h-1},\dots,\hat{p}_{t+1},\;y_{t},\dots,y_{t-d^{\prime}+1},\;C_{t}^{d_{c}},\;s_{t+h};\theta_{h}\bigr{)}.$
@@ 3.2.2 Multiple-input multiple-output (MIMO)
A multiple-output learner returns an $H$-dimensional prediction vector in a single forward pass, eliminating the need for iterative generation of successive horizons. Typical examples include neural network architectures such as multilayer perceptrons (MLP) and sequence models like LSTMs (Hochreiter and Schmidhuber, 1997; Lim and Zohren, 2021). By producing all future points simultaneously, these models can exploit cross-horizon dependencies that single-output strategies must ignore or approximate.
MIMO is the standard strategy for short-horizon multiple-output forecasting (Ben Taieb et al., 2010). A single estimator $f(\cdot;\theta)$ simultaneously returns the complete forecast vector for entire horizon $H$:
@@
3.3 Time series cross-validation (TSCV) and model evaluation
Hyperparameter tuning and model robustness assessment were performed using a sliding-window time series cross-validation (TSCV) approach (Tashman, 2000; Bergmeir and Benítez, 2012; Hyndman and Athanasopoulos, 2018). This procedure consists in chronologically splitting the data into successive training and validation windows, thereby replicating realistic forecasting conditions. It offers more reliable out-of-sample performance estimates and is particularly well-suited for time series data. In this study, TSCV is used at two key stages: first, for hyperparameter tuning based solely on the training data, as described in Algorithm 1, and second, for the final model evaluation using an independent test set. During final evaluation, the initial training window includes the full training set, and previously tuned hyperparameters are used.
Input:
  • Initial window length $m$
  • Step size $h$
  • Grid of hyperparameter configurations $\Theta$
  • Evaluation function $\operatorname{Metric}(\cdot)$
Output:  Optimal hyperparameters $\theta^{\star}$
foreach $\theta\in\Theta$ do
       $\mathcal{P}\leftarrow\varnothing$
       for $k\leftarrow 0$ to $K-1$ do
             $\mathcal{D}_{\text{train}}^{(k)}\leftarrow\{(x_{i},y_{i})\}_{i=1}^{m+kh}$;
            Train the model with $\theta$ on $\mathcal{D}_{\text{train}}^{(k)}$;
             Predict $\hat{y}_{i}$ for each $(x_{i},y_{i})\in\mathcal{D}_{\text{val}}^{(k)}$;
       end for
      $\text{score}(\theta)\leftarrow\operatorname{Metric}(\mathcal{P})$;
end foreach
Algorithm 1 Sliding-window cross-validation.
Algorithm 1 outlines the step-by-step validation process: the training window is gradually extended over time, while predictions are validated on the next window, providing a realistic short-term forecast assessment. In our study, the step size $H$ which defines the validation window length is set to 7 days to yield weekly forecasts. To address the strong class imbalance (see Table 4), we used weighted loss functions aimed at enhancing the model’s sensitivity to minority-class instances (He and Garcia, 2009).
@@ 3.4 Period-level risk assessment and evaluation metrics
While accurately predicting the exact date of an accident would be ideal, it is rarely feasible. consequently, the prevailing objective is to evaluate whether a specified time interval is characterized by elevated risk. Aggregating data by week, for instance, helps smooth out daily fluctuations and emphasizes the overall occurrence of accidents within the period. This approach is analogous to intermittent demand problems in inventory management (Croston, 1972; Syntetos and Boylan, 2005; Wallström and Segerstedt, 2010), where the focus is placed on stock availability over a period rather than on precise daily tracking.
@@
[…]
@@ 5.1 Variable overview and feature engineering
holiday categorical Indicator for the summer break
@@ 5.2 Visual assessment of stationarity
Various visualization techniques exist for binary time series (see Weiß, 2008). One of the most effective is the rate evolution graph (Ribler, 1997), shown on the right-hand side of Figure 3.
@@ 5.3 Autocorrelation and calendar effects
Figure 4 presents, in the top panel, the temporal dependence of the binary series as measured by Cohen’s Kappa statistic (Weiß and Goeb, 2008; Weiß, 2009). The autocorrelation remains weak at all lags, suggesting no strong memory effect from one day to the next.
$\boldsymbol{n_{0}}$ $\boldsymbol{p_{0}}$ $\boldsymbol{n_{1}}$ $\boldsymbol{p_{1}}$
ITW 1121 80.2% 276 19.8%
ITW-d1 1333 95.4% 64 4.6%
ExW 1258 90.1% 139 9.9%
Table 4: Frequency distribution: $n_{0}$ = number of days without accidents, $p_{0}$ = proportion of no accidents, $n_{1}$ = number of days with accidents, $p_{1}$ = proportion of accident days.
@@
5.4 Training setup and data split
Hyperparameter ITW ITW-d1 ExW
Batch size 32 32 32
Learning rate 0.01 0.01 0.01
Hidden units 32 32 32
Layers 2 2 2
Epochs 75 50 50
$dy$ 7 14 30
$dc$ 7 14 30
$\tau$ 0.4 0.6 0.5
Table 5: Optimal LSTM hyper-parameters for each time series.
The training period spans from 2019 to 2022, while data from 2022 onward were used for testing. This temporal split ensures that all evaluations are performed in a true out-of-sample setting, simulating real-world forecasting conditions. Table 6 summarizes the number of samples and the class distribution for each series across training and test sets. All three datasets (ITW, ITW-d1, and ExW) exhibit a strong class imbalance, with most days involving no accident. For instance, in the ITW-d1 test set, only 14 out of 307 days recorded an accident, i.e., less than 5%. This imbalance highlights the importance of using evaluation metrics robust to rare events. Model training and hyperparameter calibration, including the classification threshold, follow the time series cross-validation procedure described in Section 3.3. The initial training window covers 60% of the training data (665 days), and the prediction horizon is set to 7 days. The model is retrained every 7 days to ensure up-to-date forecasting. Table 5 shows the hyperparameters found for the LSTM model for the ITW, ITW-d1, and ExW series.
@@ 6 Results
Table 7 presents the performance of all methods across the three analyzed time series using four metrics: Recall (RE), Precision (PR), F1-score (F1), and balanced accuracy (BA). The naive baseline relies on a simple strategy that repeats the accident status of the same weekday from the previous week.
Model ITW ITW-d1 ExW
@@
Naive 0.65 0.67 0.66 0.37 0.15 0.15 0.15 0.39 0.58 0.58 0.58 0.51
Logistic regression 0.32 0.91 0.48 0.62 0.23 0.27 0.25 0.48 0.58 0.56 0.57 0.49
Linear discriminant analysis 0.39 0.80 0.52 0.56 0.23 0.33 0.27 0.51 0.92 0.65 0.76 0.63
Decision tree 0.23 0.78 0.35 0.52 0.00 0.00 0.00 0.50 0.67 0.64 0.65 0.58
Random forest 0.48 0.75 0.59 0.51 0.23 0.33 0.27 0.51 0.88 0.60 0.71 0.55
Histogram boosting gradient 0.13 0.57 0.21 0.43 0.15 1.00 0.27 0.58 0.58 0.64 0.61 0.57
XGBoost 0.68 0.72 0.70 0.48 0.15 0.50 0.24 0.54 0.33 0.62 0.43 0.53
LightGBM 0.48 0.64 0.33 0.43 0.08 0.50 0.13 0.52 0.67 0.57 0.62 0.50
Multi-layer perceptron 0.24 0.73 0.36 0.43 0.69 0.29 0.41 0.47 1.00 0.59 0.74 0.53
LSTM 0.77 0.87 0.81 0.63 0.85 0.79 0.82 0.87 0.79 0.70 0.75 0.67
Table 7: Performance comparison on the ITW, ITW-d1 and ExW series.
In terms of balanced accuracy, the LSTM model outperforms all other approaches on each of the three datasets. It is particularly effective at identifying both risky and safe periods, achieving a BA of 0.87 on the ITW-d1 series significantly higher than other models, which generally remain close to 0.50. Traditional machine learning models such as logistic regression, SVC, and decision trees show limited performance but superior to the naive baseline.
@@
Figure 5: Weekly accident risk prediction (top) and comparison with observed accident counts (bottom) on the ITW-d1 series with LSTM model.
Figure 5 focuses on the ITW-d1 series and visually illustrates the results obtained with the LSTM model (for similar plots related to the other series, see  C). The top plot displays the weekly accident classification. The strong overlap between observed and predicted risky periods demonstrates the model ability to capture the underlying accident patterns, although some false positives remain (predicted risky periods without actual accidents). These false alarms are generally associated with elevated risk probabilities and should not be overlooked in a prevention context. In the lower panel, the number of accidents per week is compared between observed and predicted values. Overall, the model successfully captures weekly trends in accident counts. In some weeks, the predicted number of accident days slightly exceeds the true count, but such overestimations remain reasonable and could serve as useful precautionary signals.
@@
Table 8 provides a detailed focus on three consecutive weeks from the ITW-d1 series (results for the other series are available in C). These weeks are classified respectively as risky, safe, and risky. The risky weeks contain multiple days with high predicted probabilities, supporting their classification. In contrast, the safe week is clearly identified, showing no alarming signals. This detailed view confirms the ability of our framework to dynamically capture short-term shifts in risk and to adapt to evolving safety conditions.
Week Day 1 Day 2 Day 3 Day 4 Day 5 Day 6 Day 7
3 0.996 0.620 0.002 0.001 0.012 0.001 0.013
4 0.000 0.001 0.000 0.000 0.000 0.000 0.004
5 0.107 0.999 0.088 0.000 0.008 0.001 0.000
Table 8: Daily accident probabilities for Weeks 3–5. Red cells indicate probabilities $>0.6$.

### v3

@@ 3.1 Forecasting accident risk via binary time series modeling
We represent the daily occurrence of accidents using a binary time series $\{y_{t}\}_{t=1}^{T}$, where $y_{t}=1$ if at least one accident occurs on day $t$ and $y_{t}=0$ otherwise. Two predictor families are distinguished: (i) future calendar covariates $s_{t}$ such as month or day of week, which are fully known for any future date; and (ii) dynamic inspection covariates $c_{t}$ extracted from the most recent safety inspection report available at day $t$.
@@
3.2 Multi-step forecasting strategies and evaluated machine learning models
Our aim is to predict the sequence $\{y_{t+h}\}_{h=1}^{H}$, thus producing forecasts for multiple future time steps. Several strategies can be adopted (Bontempi2013), which are commonly categorised by the output dimensionality of the underlying model. Our framework is model-agnostic. We therefore evaluate both classical single-output learners (used with the Direct–Recursive strategy) and multiple-output deep models (MIMO or encoder–decoder). Table 2 summarizes the models and their multi-step mapping.
@@ 3.2.1 Direct recursive strategy (DirRec)
A single-output learner produces one step ahead at a time. To extend it to multi-step forecasting, we use the DirRec (Direct-Recursive) strategy (Sorjamaa2006timeseries), which combines direct and recursive methods. It trains a separate estimator $f_{h}(\cdot;\theta_{h})$ for every horizon $h=1,\dots,H$. Except for the first, each estimator receives the forecasts produced at earlier horizons as additional inputs. Accordingly, the one-step-ahead forecast is
$\hat{p}_{t+1}\;=\;f_{1}\!\bigl(y_{t},\dots,y_{t-d_{y}+1},\;C_{t}^{d_{c}},\;s_{t+1};\theta_{1}\bigr).$
@@
$\hat{p}_{t+h}\;=\;f_{h}\!\bigl(\hat{p}_{t+h-1},\dots,\hat{p}_{t+1},\;y_{t},\dots,y_{t-d_{y}+1},\;C_{t}^{d_{c}},\;s_{t+h};\theta_{h}\bigr).$
@@ 3.2.2 Multiple-input multiple-output (MIMO)
A multiple-output learner returns an $H$-dimensional prediction vector in a single forward pass, eliminating the need for iterative generation of successive horizons. By producing all future points simultaneously, these learners can exploit cross-horizon dependencies that single-output strategies must ignore or approximate.
MIMO is the standard strategy for multiple-output forecasting (BenTaieb). A single estimator $f(\cdot;\theta)$ simultaneously returns the complete forecast vector for entire horizon $H$:
@@
3.2.3 Autoregressive Seq2Seq architecture
We also consider an encoder–decoder (Seq2Seq) architecture with an autoregressive decoder (Sutskever2014). An encoder maps the recent history $\big(Y_{t}^{d_{y}},C_{t}^{d_{c}}\big)$ into a latent state
$\mathbf{z}_{t}\;=\;\mathrm{enc}\!\bigl(Y_{t}^{d_{y}},C_{t}^{d_{c}};\,\psi\bigr),$
where $\psi$ are the encoder parameters. and a decoder then produces accident probabilities step by step, each conditioned on the previous output and on the known future calendar covariates:
$\hat{p}_{t+h}\;=\;\sigma\!\Big(\mathrm{dec}\!\big(\mathbf{z}_{t},\;u_{t+h-1},\;s_{t+h};\,\phi\big)\Big),\qquad h=1,\dots,H,$
with $u_{t}=y_{t}$ (last observed outcome) for $h=1$ and, for $h\geq 2$, $u_{t+h-1}=\hat{p}_{t+h-1}$. Moreover, $\sigma(\cdot)$ denotes the logistic function and $\phi$ are the decoder parameters. While MIMO predicts the entire vector $(\hat{p}_{t+1},\dots,\hat{p}_{t+H})$ in a single forward pass with a shared parameter set, an autoregressive seq2seq decoder predicts one step at a time, each step using the previous output. This makes the forecast length flexible (the decoder can be unrolled to any $H$) and lets day-to-day effects carry over, but early errors may propagate to later steps. By contrast, MIMO avoids this roll-out effect but captures cross-horizon structure only implicitly through its shared representation. A non-autoregressive encoder–decoder can also produce all horizons in a single forward pass. it avoids roll-out error but does not condition on previous predictions.
Family Model Output type Multi-step strategy References
Machine learning Logistic regression Single-output DirRec (bishop2006pattern)
Linear discriminant analysis Single-output DirRec (bishop2006pattern)
Decision tree (CART) Single-output DirRec (breiman1984cart)
Random forest Single-output DirRec (breiman2001random)
HistGradient Boosting Single-output DirRec (friedman2001greedy; ke2017lightgbm)
XGBoost Single-output DirRec (xgboost; friedman2001greedy)
LightGBM Single-output DirRec (ke2017lightgbm)
Deep learning Multilayer perceptron (MLP) Vector of size $H$ MIMO (bishop1995nn)
LSTM–MIMO Vector of size $H$ MIMO (HochreiterSchmidhuber1997; taieb2015review)
LSTM–Seq2Seq One-step output Seq2Seq (autoregressive) (Sutskever2014; cho2014seq2seq)
TCN Vector of size $H$ MIMO (bai2018tcn)
TFT Vector of size $H$ Seq2Seq (non-autoregressive) (Lim2021TFT)
Table 2: evaluated models, output types and multi-step strategies.
@@ 3.3 Period-level risk assessment and evaluation metrics
While accurately predicting the exact date of an accident would be ideal, it is rarely feasible. consequently, the prevailing objective is to evaluate whether a specified time interval is characterized by elevated risk. Aggregating data by week, for instance, helps smooth out daily fluctuations and emphasizes the overall occurrence of accidents within the period. This approach is analogous to intermittent demand problems in inventory management (Croston1972; SyntetosBoylanCroston2005; wallstrom2010evaluation), where the focus is placed on stock availability over a period rather than on precise daily tracking.
@@
$\text{F1}=2\,\cdot\,\frac{\text{PR}\cdot\text{RE}}{\text{PR}+\text{RE}}.$
@@ 5.1 Variable overview and feature engineering
weekend_day categorical Weekend indicator (saturday or sunday)
holiday categorical Indicator for the summer break and public holidays
@@ 5.2 Visual assessment of stationarity
Various visualization techniques exist for binary time series (see weiss2008visual). One of the most effective is the rate evolution graph (ribler1997visualizing), shown on the right-hand side of Figure 3.
@@ 5.3 Autocorrelation and calendar effects
Figure 4 presents, in the top panel, the temporal dependence of the binary series as measured by Cohen’s Kappa statistic (WeissGoeb2008; Weiss2009). The autocorrelation remains weak at all lags, suggesting no strong memory effect from one day to the next.
@@
6 Training setup and evaluation
The training period spans from 2019 to 2022, while data from 2022 onward were used for testing. This temporal split ensures that all evaluations are performed in a true out-of-sample setting, simulating real-world forecasting conditions. Table 5 summarizes the number of samples and the class distribution for each series across training and test sets. All three datasets (ITW, ITW-d1, and ExW) exhibit a strong class imbalance, with most days involving no accident. For instance, in the ITW-d1 test set, only 14 out of 289 days recorded an accident, i.e., less than 5%. This imbalance highlights the importance of using evaluation metrics robust to rare events and using class-weighted loss functions to improve sensitivity to the minority class (He2009Imbalanced). Model training and hyperparameter calibration, including the classification threshold, follow the time series cross-validation procedure described in Algorithm 1. The initial training window covers 60% of the training data, and the prediction horizon is set to 7 days. The model is retrained every 7 days to ensure up-to-date forecasting.For reproducibility, D reports the hyperparameter search spaces and the selected operating points for all models considered in this study for the ITW, ITW-d1, and ExW.
@@ 6 Training setup and evaluation
Hyperparameter tuning and model robustness assessment were performed using a sliding-window time series cross-validation (TSCV) approach (Tashman2000; bergmeir2012crossval; hyndman2018forecasting). This procedure consists in chronologically splitting the data into successive training and validation windows, thereby replicating realistic forecasting conditions. It offers more reliable out-of-sample performance estimates and is particularly well-suited for time series data. In this study, TSCV is used at two key stages: first, for hyperparameter tuning based solely on the training data, as described in Algorithm 1, and second, for the final model evaluation using an independent test set. During final evaluation, the initial training window includes the full training set, and previously tuned hyperparameters are used.
Input:
  • Initial window length $m$
  • Step size $h$
  • Grid of hyperparameter configurations $\Theta$
  • Evaluation function $\operatorname{Metric}(\cdot)$
Output:  Optimal hyperparameters $\theta^{\star}$
foreach $\theta\in\Theta$ do
 $\mathcal{P}\leftarrow\varnothing$
 for $k\leftarrow 0$ to K1K-1 do
    $\mathcal{D}_{\text{train}}^{(k)}\leftarrow\{(x_{i},y_{i})\}_{i=1}^{m+kh}$;
     Train the model with $\theta$ on $\mathcal{D}_{\text{train}}^{(k)}$;
      Predict $\hat{y}_{i}$ for each $(x_{i},y_{i})\in\mathcal{D}_{\text{val}}^{(k)}$;
   end for
 $\text{score}(\theta)\leftarrow\operatorname{Metric}(\mathcal{P})$;
end foreach
Algorithm 1 Sliding‑window cross‑validation.
Algorithm 1 outlines the step-by-step validation process: the training window is gradually extended over time, while predictions are validated on the next window, providing a realistic short-term forecast assessment. In our study, the step size $H$ which defines the validation window length is set to 7 days to yield weekly forecasts.
@@ 7 Results
7.1 Operational baselines
To anchor the evaluation in practical routines, we benchmarked the learning models against two simple, operations-focused, training-free baselines reflecting rules commonly used by safety teams.
7.1.1 Rolling accident frequency over $W$ days
Following standard safety monitoring practices, the rolling frequency baseline approximates the short-term accident rate over the last $W$ days. This design is conceptually aligned with the total recordable incident rate (TRIR), a well-established safety indicator that aggregates recent event counts over fixed time windows to track variations in safety performance (hse2001measuring). Such trailing indicators are consistent with operational safety management approaches, where repeated minor incidents or a rise in local accident frequency are regarded as early warning signals of deteriorating safety conditions (goh2012organizational; reason1997managing).
A trailing accident rate is computed
$r_{t}=\frac{1}{W}\sum_{i=0}^{W-1}y_{t-i},$
and we use it as a constant risk score for the whole upcoming week:
$\hat{p}_{t+h}=r_{t},\qquad h=1,\dots,7.$
To convert this score into a period-level alert, we flag the upcoming week as at risk whenever $r_{t}\geq\tau$. The threshold $\tau$ and the window length $W$ (chosen from $\{7,14,28\}$) are selected on the training split by maximizing period-level balanced accuracy, and then held fixed for evaluation.This procedure yields W=14W{=}14 and thresholds $\tau_{\text{ITW}}{=}0.4$, $\tau_{\text{ITW-d1}}{=}0.4$, and $\tau_{\text{EXW}}{=}0.5$.
7.1.2 Naive
This baseline copies last week’s pattern into the next week:
$\hat{y}_{t+h}=y_{t+h-7},\qquad h=1,\dots,7.$
It reflects the intuition that short-term risk often repeats on the same weekday.
7.2 Model performance assessment
Table 6 compares all methods on the three series using Recall (RE), Precision (PR), F1, and balanced accuracy (BA). Across datasets, learned models consistently outperform the operational baselines, confirming the usefulness of the proposed framework. On ITW and ITW-d1, the LSTM (MIMO) achieves the highest BA, indicating superior ability to detect risky weeks while limiting false alarms. On ExW, a simple Decision Tree attains the best BA, suggesting that a low-capacity, rule-like boundary fits that series’ patterns particularly well. Tree/boosting models and the MLP are competitive but generally below the best deep sequence model on ITW and ITW-d1. The rolling-frequency baseline (Freq. rolling) remains informative yet is surpassed by most learned models, which better exploit temporal dependencies and calendar effects. Across the three series, LSTM–Seq2seq, TCN, and TFT are consistently competitive. On ITW and ITW-d1, they generally rank below the top LSTM–MIMO yet clearly above the operational baselines, and they are broadly comparable to the stronger tree/boosting and MLP models. On ExW, none of the three sequence learners takes the lead; the best score is obtained by a Decision Tree, while the sequence models remain competitive.
Model family Model ITW ITW-d1 ExW
@@ 7.2 Model performance assessment
Operational baselines Naive (last week) 0.66 0.68 0.67 0.37 0.15 0.15 0.15 0.39 0.56 0.58 0.57 0.50
Freq. rolling (W) 0.69 0.79 0.73 0.57 0.08 0.33 0.12 0.50 0.52 0.65 0.58 0.56
Machine learning Logistic regression 0.37 1.00 0.54 0.69 0.46 0.75 0.57 0.70 0.78 0.67 0.72 0.67
Linear discriminant analysis 0.59 0.79 0.68 0.57 0.50 0.29 0.36 0.51 0.88 0.66 0.76 0.63
Decision tree 0.81 0.87 0.84 0.72 0.46 0.66 0.54 0.68 0.64 0.84 0.73 0.74
Random forest 0.62 0.82 0.71 0.66 0.92 0.38 0.54 0.64 0.88 0.65 0.74 0.61
Histogram boosting gradient 0.93 0.81 0.87 0.65 0.38 0.71 0.50 0.66 0.28 0.87 0.42 0.61
XGBoost 0.69 0.88 0.77 0.71 0.85 0.46 0.59 0.71 0.48 0.80 0.60 0.67
LightGBM 0.78 0.86 0.81 0.70 0.92 0.43 0.58 0.69 0.76 0.65 0.70 0.60
Deep learning Multi-layer perceptron 0.62 0.90 0.74 0.72 0.69 0.47 0.56 0.68 0.64 0.76 0.69 0.68
LSTM – MIMO 0.71 0.92 0.80 0.73 0.85 0.79 0.82 0.87 0.79 0.70 0.75 0.67
Seq2seq 0.69 0.88 0.77 0.70 0.53 0.70 0.60 0.72 0.64 0.69 0.66 0.62
TCN 0.65 0.91 0.76 0.71 0.92 0.50 0.51 0.76 0.52 0.76 0.62 0.65
TFT 0.62 0.83 0.71 0.63 0.91 0.39 0.55 0.68 0.13 1.00 0.23 0.56
Table 6: Performance comparison on the ITW, ITW-d1 and ExW series on the test set.
@@
Figure 5: Weekly accident risk prediction (top) and comparison with observed accident counts (bottom) on the ITW-d1 series with LSTM-MIMO model.
Figure 5 illustrates these findings at the weekly level. The top panel shows a close alignment between observed and predicted risky weeks; remaining false positives occur when the model assigns high probabilities close to the decision threshold, which can still be valuable as preventive alerts. The bottom panel compares weekly accident counts and shows that the model tracks week-to-week variations reasonably well, with mild overestimation in a few weeks that remains acceptable in a prevention setting.
Figure 6: Daily accident probabilities for Weeks 21,22 and 23. Red cells indicate probabilities >0.6>0.6
@@
Table 6 provides a detailed focus on three consecutive weeks from the ITW-d1 series (results for the other series are available in C). These weeks are classified respectively as safe, risky, and safe. The risky week contain one day with high predicted probability, supporting their classification. In contrast, the safe week is clearly identified, showing no alarming signals. This detailed view confirms the ability of our framework to dynamically capture short-term shifts in risk and to adapt to evolving safety conditions.
7.3 Horizon sensitivity and calendar effects
Here, We assess how the decision period length or horizon ($h\in\{3,5,7\}$) impacts period-level performance and how known future calendar covariates $s_{t+h}$ impact the performance of our models.
7.3.1 Horizon sensitivity
Figure 7 shows that balanced accuracy improves with longer horizons, peaking at the weekly horizon h=7\,h=7\, across all series. This pattern can be explained by two complementary effects: aggregating decisions over seven days smooths the noise inherent to rare events and yields more stable period-level signals, and several operational drivers (for example, inspection cadence and work scheduling) naturally follow weekly rhythms. The largest improvement is observed for ITW-d1, where events are concentrated in a single department and weekly aggregation therefore provides particularly informative signals.
From an operational perspective, a one-week look-ahead constitutes a natural planning unit: it affords sufficient lead time to allocate resources (e.g., targeted safety briefings or housekeeping actions), schedule interventions, and communicate priorities, while remaining specific enough to be actionable. Shorter horizons ($h=3$ or $h=5$) can be more timely but tend to be noisier and less aligned with organisational processes, which explains their lower period-level balanced accuracy.
Figure 7: Period-level Balanced accuracy by forecast Horizon (Best model per series).
7.3.2 Calendar ablation at weekly horizon.
To isolate the marginal contribution of calendar features at h=7h{=}7, we rerun the identical backtesting protocol after removing $s_{t+h}$ from the future covariates, keeping all else fixed (same folds, same training data, same optimization settings, and the operating threshold $\tau$ fixed to the value calibrated in the full model). Let $\mathrm{BA}^{(\text{cal})}(D)$ denote the period-level balanced accuracy on dataset $D$ with calendar features, and $\mathrm{BA}^{(\neg\text{cal})}(D)$ the BA without them. We report the absolute gain
$\Delta\mathrm{BA}(D)\;=\;\mathrm{BA}^{(\text{cal})}(D)-\mathrm{BA}^{(\neg\text{cal})}(D).$
[…]
ITW LSTM–MIMO 0.71 0.59 ++0.12
ITW-d1 LSTM–MIMO 0.87 0.61 ++0.26
ExW Decision tree 0.74 0.64 ++0.10
Table 7: Period-level balanced accuracy (BA) at h=7h{=}7 for the best model of each series (selected according to BA), with and without future calendar covariates.
Consistent, positive $\Delta\mathrm{BA}$ across all series indicates that weekday/weekend and holiday structure is informative at the weekly scale (Table 7). The gain is lower for ITW (+0.12+0.12) and EXW (+0.10+0.10), suggesting that calendar signals are diluted when aggregating multiple departments with heterogeneous rhythms and that short-term history already captures weak weekly recurrences. In contrast, ITW-d1 shows a large improvement (+0.26+0.26), consistent with a more regular weekly pattern within a single department where calendar cues are highly predictive.